OpenStudy (anonymous):

Which postulate can be used to prove the two triangles below are similar? Explain your answer using complete sentences, and provide evidence to support your claims.

OpenStudy (anonymous):

I don't know what postulate they are similar by. These triangles are similar by the . Each of the triangle's angles have to be equal to 180 degrees. We should add all the angles up and see if they equal 180 degrees. 50+33+97 = 180 Since the sum of the angles are 180 degrees, we can conclude that the two triangles's angles are similar by corresponding angles.

OpenStudy (epoweritheta):

no no any triangle has sum of angles 180 we cant decide by that

OpenStudy (epoweritheta):

what is angle C in the first figure ??

OpenStudy (epoweritheta):

and B' in the second figure ??

OpenStudy (anonymous):

Angle c in the first figure is equal to 97 degrees. And B' in the second is equal to 33 degrees.

OpenStudy (epoweritheta):

good since all the angles of these triangles are equal we have by AAA postulate the triangles are similar . got it ???

OpenStudy (anonymous):

So I can say "If you calculate the missing angle to equal 180 for both triangles, and they are equal, they are similar by the AAA postulate?

OpenStudy (epoweritheta):

no no if all the angles are equal , you can use AAA

OpenStudy (epoweritheta):

see "A School Geometry " a free book that teaches geometry from basic

OpenStudy (anonymous):

These triangles are similar by the AAA postulate. Each of the triangle's angles have to be equal to 180 degrees. We should find the missing angles for the triangles. For Triangle A: 50+33+? = 180 50+33+97 = 180 For Triangle B: 50+97+? = 180 50+97+33 = 180 Since the sum of the angles are 180 degrees, we can conclude that the two triangles's angles are similar by AAA.

OpenStudy (anonymous):

Or am I wrong? (I will check the book in a minute.)

OpenStudy (epoweritheta):

no no 50=50 ,33=33,97=97 so AAA okay ; not 180 ; any triangle angles sum to 180 but may not be similar

OpenStudy (anonymous):

So: These triangles are similar by the AAA postulate. Each of the triangle's angles have to be equal to 180 degrees. We should find the missing angles for the triangles. For Triangle A: 50+33+? = 180 50+33+97 = 180 For Triangle B: 50+97+? = 180 50+97+33 = 180 If 50-50, 33=33, and 97=97, we can conclude that the triangles are similar by the AAA postulate.

OpenStudy (anonymous):

The 50-50 was supposed to be 50=50.

OpenStudy (anonymous):

Is it correct?

OpenStudy (epoweritheta):

ya ya

OpenStudy (anonymous):

Thanks! :D

OpenStudy (epoweritheta):

;)